AAVSOnet news 2013-01-25
SRO50 and BSM-South continue working reliably. Here is an update on some other telescopes in the network!
BSM_CA - this Bright Star Monitor at Bob Stine's observatory in California is nearly ready to go. John Gross is getting the autofocusing to work and setting up ACP/ACP Scheduler. Bob is at his "Ranchito" about half time, with the next visit in a few days, so we are hoping to have this system taking survey data in early February.
BSM_Berry - this system is on top of HQ. I just finished polar aligning the system, and have a small amount of cable rerouting to do. The faint limit is fine for BSM targets, even in the heart of Boston! I expect it to be taking survey data in early February.
OC61 - the main hardware glitch that caused the telescope to run away has been identified. The University of Canterbury instrument shop has made some improvements in preparation for the installation of the spectrograph. These should be added to the system next week, with some shakedown time required. Otherwise, it is ready to go and will be taking data within a few days.
TMO61 - the RA drive is not running as smoothly as Jon Holtzman would like, so he is making some adjustments. I've ordered the final QSI-683wsg science camera. The system has already taken several nights of automated data, so we expect it to be fully operational within days.
W30 - Bill Goff has this system up and running, with one night of automated operation completed. As soon as the bad weather moves away from California, we will be updating the queues and putting it into full service.
W28, BSM_NM will be installed in a few months when the weather gets warmer. I will be on vacation next week, which is why many of the systems are due to be finished the week after that.